The chance is 0.35
Is total area under a normal distribution finite?
Yes, and is equal to 1. This is true for normal distribution using any mean and variance.
What is the probabilty of not getting a 9 on two dice?
There are 36 total ways two dice can be rolled. 4 of those give you a 9. The probability is 4/36=1/9
What is the probability of drawing two consecutive aces from a deck of fifty-two cards?
It is 1/221.
Assume that the standard deck is completely shuffled in a completely unbiased way.
The probability of drawing the first ace is 4/52, since there are 4 aces in a standard deck.
The probability of drawing the second ace is 3/51, since there are three aces remaining and 51 cards from which to choose.
12/52 X 51 equals 12/2652, which equals 1/221.
Why is it called experimental probability?
It's difficult to think of a real event to which an exact probability can be assigned. We say that flipping a coin yields 'heads' with probability 1/2 but we do not know that definitely. The only way of assigning a probability in the sense of numbers of heads versus total numbers of flips is by experiment. (Be aware though that there are other interpretations of the word probability.) If I were to flip a coin 500 times and obtained 249 heads then the experimental probability of obtaining a head would be 249/500 or 0.498.
Wherever there is some random (that is, non-determinate) element in the occurrence of an event.
What is the first step in finding the probability of getting two numbers whose sum 2?
The first step is to determine what the experiment is: rolling a pair of dice, drawing two cards from a deck of playing cards, selecting numbered cards?
How many even numbers in a regular deck of cards?
there's 4 suits in one deck of cards.
if j =11, queen =12, and king =13.
that means there will be 6 even numbers card in one suit, and times that by 4, which mean 24. theres 24 cards are even in a regular deck of cards.
What is the probability of having a heart if you have 8 cards from a deck?
Prob(Having a heart) = 1 - Prob(No hearts)
= 1 - (39/52)*(38/51)*(37/50)*(36/49)*(35/48)*(34/47)*(33/46)*(32/45)
= 1 - 2480637519360/30342338208000
= 1 - 0.0818 = 0.9182
no equal answer
What is the probability the cube will land on 4 and the coin will land heads up?
There is a 1/6 chance of rolling a 4 on a fair die, and a 1/2 chance of a fair coin landing heads up. Multiply 1/6 X 1/2. The probability of both happening is 1/12.
What is a probability of rollong an even number on a number diece?
A die has 6 sides with the numbers 1,2,3,4,5 and 6. Three of those numbers are even. So the probability of rolling an even number is 3 out of 6, or 1 out of 2.
Probability of hitting the bull's-eye on the next random shot after eliminating . . .
- no answer . . . . . 20%
- 1 answer . . . . . . 25%
- 2 answers . . . . . 331/3%
- 3 answers . . . . . 50%
- 4 answers . . . . . 100%
- 5 answers . . . . . zero
Probability of choosing the one correct answer increases significantly from
these figures if you actually know something about the subject.
Because you can't live with out them not being married to you.
What is a frequency distribution defined by its average and standard deviation?
The triangular, uniform, binomial, Poisson, geometric, exponential and Gaussian distributions are some that can be so defined. In fact, the Poisson and exponential need only the mean.
A single die is rolled what is the probability tha the number rolled will be greater than 3?
Total different outcomes = 6
Successful outcomes = 3 (rolls of 4, 5, or 6)
Probability of success = 3/6 = 1/2 = 50%
8 dices are thrown probability of even numbers?
Since there are only odd & even numbers, the probability is 1/2 or 0.5.
What is the probability of drawing an even number or an ace?
If enough cards are drawn it is a certainty. In a single draw, the probability is 0.3846 if you assume that Q is neither odd nor even. If you take Q to be 12, and therefore, even, then the probability increases to 0.4615, approx.
What are the benefits of a two tailed test?
They are used to test hypothesis such as the mean is some value where you do not know if otherwise the mean is less or more.
What does the probability density function tell you?
The area under the pdf between two values is the probability that the random variable lies between those two values.